Find the speed of an ocean wave whose displacement is given by the equation y=3.7cos(2.2x-5.6t) where x and y are in meters and t is in seconds?

A) 2.5m/s

B) 3.5m/s

C) 1.9m/s

D) 4.5m/s
Answer #1

here k = 2pi/L   = 2*3.14/L = 2.2

L = 2*3.14/2.2

L = 2.85 m

again W = 2pif = 5.6

so f = 5.6/.2pi

f = 0.891 Hz

so

V = Lf

so

V = 2.85 * 0.891

V = 2.5 m/s

option A it is
Answer #2

y = a cos(kx -?t+?0) = a sin(kx-?t +?1), where?1= ?0+? /2. Here?0= 0
Answer #4

I think that D